# Fully Dynamic Euclidean $k$-Means

### v2

@@ 5.1 Preliminaries
Throughout this section, we assume that the input space to our algorithm is $[\Delta]^{d}$, where the dimension $d$ is $O(\log n)$. This assumption can be achieved by the well-known Johnson-Lindenstrauss transformation [JL84]. Assume the input space is $\mathbb{R}^{m}$. We initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$. Whenever a point $x\in\mathbb{R}^{m}$ is inserted to the dynamic space, we first apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $[\Delta]^{d}$. As a result, we can feed this new dynamic space into our algorithm. Since $d=O(\log n)$, we conclude that the running time performance of all the data structures in Section 4 becomes $\tilde{O}(n^{O(\epsilon)})$. Then, by scaling $\epsilon$, we assume that the running time of these data structures is $\tilde{O}(n^{\epsilon})$. For simplicity, we consider a parameter $p_{\epsilon}$ which is a polynomial of $(1/\epsilon)$ and is an upper bound on the approximation guarantees of all data structures in Section 4. So, we consider the following throughout the entire Section 5.
@@
approximation guarantee at most $\text{approximation guarantee at most }p_{\epsilon}\text{ and running time at most }\tilde{O}(n^{\epsilon}).$
@@
Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $[\Delta]^{d}$. We call this sequence $t$-robust w.r.t. the current data set $X$, if there exists a sequence $(B_{0},B_{1},\cdots,B_{t})$ of subsets of $X$ such that the following hold,
@@
$u\text{ is at least }t\text{-robust, where }t\text{ is the smallest integer satisfying }p_{\epsilon}^{6t}\geq\mathrm{dist}(u,S-u)/p_{\epsilon}^{20}.$ (9)
@@ 5.2 Implementation
$i=0$to $\log_{2}k$
[…]
After finding $\widehat{\ell}$, we set the length of the epoch to be $\ell+1$ where ^$\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot p_{\epsilon}^{102}}\right\rfloor$.
@@ 5.3 Implementing Robustify
$t[u]\geq t,\text{ where }t\text{ is the smallest integer satisfying }p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{20}$ (13)
@@ 5.4 Searching for Non-Robust Centers
The proof of this claim follows from [BCF25a], but for the sake of completeness, we provide the proof in Part III (see Section 20.10).
@@ 5.4.1 Implementing Part 1
This claim shows that every update in the data set can contaminate at most one center in each scale $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$. The proof of this claim follows from [BCF25a], but for the sake of completeness, we provide the proof in Part III (see Section 20.11).
@@ 5.4.2 Implementing Part 2
This lemma follows from combining the ANN and [BGJ+24, Theorem 3.1]. We have already implemented ANN in Section 4, and in [BGJ+24], it is explained how we can implement the above data structure by using an ANN.
@@ 5.4.3 Termination of Robustify
After each change in the center set $S$, the distances of centers $u\in S$ to other centers might change. This means that new centers might be inserted into $\mathcal{Y}$ and they might need a subsequent call to MakeRobust. Similar to [BCF25a], it is possible to show that if a call MakeRobust is made on a center in Robustify, it will continue to satisfy Conditions (13) and (9) until the end of the call to the current Robustify, and our procedure will not make another call on that center.
@@
Consider any call to Robustify$(W^{\prime})$, and suppose that it sets $w_{0}\leftarrow\texttt{MakeRobust}(w)$ during some iteration of the while loop. Then in subsequent iterations of the while loop in the same call to Robustify$(W^{\prime})$, we will not make any call to MakeRobust$(w_{0})$.
The proof of this lemma follows from [BCF25a] (although we have new definitions), and we provide the proof in Part III for the sake of completeness (see Section 20.9).
@@
[…]
6:$u\leftarrow$pop$(\mathcal{Y})$
[…]
@@ 5.5 Make-Robust
In this section, we introduce the subroutine MakeRobust that given $u\in S$, swaps this center with a close center $v$ which is robust (i.e. satisfies Condition Equation 13 as well as Equation 9). We start by computing the smallest integer $t$ satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{14}$. The goal is to follow the Definition 5.1. We define $x_{t}:=u$ and by iterating over $j=t$ down to 00, we find $x_{j-1}$ and $B_{j}$ satisfying the properties in Definition 5.1. We use Lemma 4.5, and make a query on this data structure for $x=x_{j}$ and $\rho=p_{\epsilon}^{6j}$. It returns a center $c^{\star}\in[\Delta]^{d}$, together with numbers $\widehat{\textsc{cost}}_{c^{\star}}$, $\widehat{\textsc{cost}}_{x_{j}}$, and $b_{x_{j}}$.
@@
[…]
3:$x_{j-1}\leftarrow x_{j}$Else
[…]
5:$v\leftarrow x_{0}$
6:$S\leftarrow S-u+v$
7:Save $t[v]:=t$ together with $v$
@@
The analysis of the approximation ratio of the algorithm follows from the prior work [BCF25a] with deliberate changes in the parameters inside the proofs. There are only two important considerations as follows. First, our new definition of robust centers still satisfy the main properties of robust centers as in [BCF25a] (see Section 15.1.1). Second, the analysis of [BCF25a] depends on the fact that the integrality gap of the standard LP-relaxation for $k$-median is constant. Since the same fact holds for $k$-means (see Theorem 1.1 in [ANFSW20]), we can extend the analysis for our purpose. For the sake of completeness, we provide a thorough analysis in Part III.
@@
The recourse analysis of the algorithm also follows from the previous work [BCF25a] with deliberate changes. There are only some considerations as follows. First, in augmented $k$-means (Lemma 4.9), instead of finding a set $A$ of size exactly $a$, we find a set of size $\tilde{O}(a)$. This increases the recourse of the algorithm by some $\log$ factors, but the same analysis works as in [BCF25a]. Second, we changed the Condition (9) to a more restricted Condition (13), and we might call MakeRobust more often. But, the same analysis as in [BCF25a] can be applied here to show that the total amortized recourse of Robustify is still $\tilde{O}(1)$. For the sake of completeness, we provide a thorough analysis in Part III.
@@ 6.1.3 Running Time of Robustify
For each of these centers $u\in\mathcal{Y}$, during 6 to 7, the algorithm decides in $\tilde{O}(1)$ time whether or not a call to MakeRobust will be made on $u$ [^7] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 4.6. As a result, the amortized time spend on part 2 is $\tilde{O}(n^{\epsilon})$.
@@ 6.1.4 Running Time of MakeRobust
The main for loop in the implementation consists of at most $O(\log_{p_{\epsilon}^{2}}(\sqrt{d}\Delta))$ many iterations. Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 4.5) used in 2 of Algorithm 3. It is obvious that all other parts of the implementation take at most $\tilde{O}(1)$ time. Hence, the total time spent to perform one call to MakeRobust is at most $\tilde{O}(n^{\epsilon})$.
@@ 6.2.2 Correctness of MakeRobust
In this part of the paper, we provide several dynamic data structures that are useful in subsequent parts. We begin by summarizing the formal statement of these data structures in Section 7. As a crucial foundation of our data structures, we design a time-efficient consistent hashing scheme in Section 8. Then, we elaborate on the construction and analysis of each data structure in Sections 9, 10, 11 and 12.
@@
Throughout this part, we fix a small constant $10<\epsilon<$ 1, denote by $X\subseteq[\Delta]^{d}$ and $w:X\to\mathbb{R}_{\geq 0}$ a dynamic weighted point set and its weight function, respectively, and denote by $S\subseteq[\Delta]^{d}$ a dynamic (unweighted) center set. Without loss of generality, we assume that both $X$ and $S$ are initially empty. We use $\tilde{O}(\cdot)$ to hide the $\mathrm{poly}(d\epsilon^{-1}\log(\Delta))$ factors.
@@ 7 Main Data Structures
The following Lemmas 7.1, 7.2, 7.3 and 7.4 are established in Section 9, where we design a unified range query structure for handling a class of “mergeable” query problems. Specifically, Lemmas 7.1, 7.2 and 7.3 address a set of approximate nearest neighbor (ANN) problems, while Lemma 7.4 concerns the $1$-means problem over approximate balls. Here, an “approximate ball” refers to a subset that is sandwiched between two balls with the same center and (up to a constant) roughly the same radii.
@@
The following Lemmas 7.5 and 7.6 address two crucial subproblems – restricted $k$-means and augmented $k$-means – respectively, which play a central role in our dynamic algorithm described in Part III. We present the data structures of Lemmas 7.5 and 7.6 in Sections 11 and 12, respectively, both of which are built on a common data structure that (implicitly) maintains an approximate assignment from the point set $X$ to the center set $S$, as described in Section 10.
@@
$\textsc{cost}(X,S-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
@@
$\textsc{cost}(X,S+A)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(1)\cdot\min_{A^{\prime}\subseteq[\Delta]^{d}:|A^{\prime}|=a}\textsc{cost}(X,S+A^{\prime}).$
@@ 8 Efficient Consistent Hashing
    (efficiency) For every point $x\in[\Delta]^{d}$, both the hash value $\varphi(x)$ and a size-$(\leq\Lambda)$ subset $\Phi_{x}\subseteq\Phi$ satisfying $\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho)$ can be computed in time $\tilde{O}(\Lambda)$.
@@
For any $\Gamma\geq 2\sqrt{2\pi}$, $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, and $\rho>0$, there exists a random hash $\varphi$ on $[\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
@@
Our construction of an efficient consistent hash builds on a previous construction from [FJL+25], which provides a weaker guarantee of consistency that holds in expectation. In addition, we leverage the special structure of their hash function – based on randomly shifted grids – and show that the set of hash values over any given ball can be computed efficiently (see Item 4 below). We summarize these results in the following lemma.
Lemma 8.3 (Hashing with Weaker Consistency Guarantees [FJL+25]).
@@
Items 1, 2 and 3 are explicitly stated in [FJL+25]. We now prove Item 4.
We review the construction from [FJL+25]: Let $v\sim\operatorname{Unif}([0,r/\sqrt{d}]^{d})$ be a uniformly random real vector. The hash function is defined as $\tilde{\varphi}:x\mapsto\lfloor\frac{x+v}{r/\sqrt{d}}\rfloor$ for $x\in[\Delta]^{d}$. Herein, the integer vector $\lfloor z\rfloor:=(\lfloor z_{1}\rfloor,\dots,\lfloor z_{d}\rfloor)\in\mathbb{Z}^{d}$ coordinatewise rounds down a real vector $z=(z_{1},\dots,z_{d})\in\mathbb{R}^{d}$.
@@
$\widetilde{\varphi}^{-1}(z)$ $\textstyle\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\big{[}z_{j}\cdot r/\sqrt{d}-v_{j},(z_{j}+1)\cdot r/\sqrt{d}-v_{j}\big{)},$
$\widetilde{\varphi}^{-1}(z^{\prime})$ $\textstyle\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\big{[}z^{\prime}_{j}\cdot r/\sqrt{d}-v_{j},(z^{\prime}_{j}+1)\cdot r/\sqrt{d}-v_{j}\big{)}.$
@@
  • $\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\leq\rho$, for every hash value $z\in\mathbb{Z}^{d}$. $\triangleright$ Item 1 of Lemma 8.3
  • $\mathbb{E}\big{[}|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\big{]}\leq\frac{\Lambda}{2C}$, for every point $x\in[\Delta]^{d}$. $\triangleright$ Item 2 of Lemma 8.3
@@
$\varphi(x)\leavevmode\nobreak\ :=\leavevmode\nobreak\ (c_{x},\widetilde{\varphi}_{c_{x}}(x)),$ $\forall x\in[\Delta]^{d}.$ (16)
@@
First, the diameter bound and the consistency properties (Items 1 and 2 of Definition 8.1) follow from Claims 8.5 and 8.6, respectively.
@@
By construction, our consistent hash function $\varphi(x)=(c_{x},\widetilde{\varphi}_{c_{x}}(x))$ induces a natural partition of the input space $[\Delta]^{d}=\cup_{c\in[C]}G_{c}$, where $G_{c}:=\{x\in[\Delta]^{d}\mid c_{x}=c\}$, $\forall c\in[C]$. In this manner, we have $|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\sum_{c\in[C]}|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c})|$, so it suffices to show that
@@
Second, the efficiency guarantee (Item 3 of Definition 8.1) directly follows from a combination of Claims 8.7 and 8.8.
@@
Following Lemma 8.3 and the stop condition “$|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, the depth-first search for every single $c=1,2,\dots,C$ stops in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$.
@@
Following Lemma 8.3 and the stop condition “$|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, every iteration $c=1,2,\dots,C$ will stop in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$ and augment the subset $\Phi_{x}$ by at most $\frac{\Lambda}{C}$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$ and the total size $|\Phi_{x}|\leq\Lambda$.
@@
(ii) $z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$, given our construction of the hash function $\varphi$ (Equations 15 and 16).
@@
[…]

### v3

@@ 6.1 Preliminaries
Throughout this section, we assume that the input space to our algorithm is $\Delta]^{d}$, where the dimension $d$ is $O(\log n)$. This assumption can be achieved by the well-known Johnson-Lindenstrauss transformation [JL84]. Assume the input space is $\mathbb{R}^{m}$. We initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$. Whenever a point $x\in\mathbb{R}^{m}$ is inserted to the dynamic space, we first apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $\Delta]^{d}$. As a result, we can feed this new dynamic space into our algorithm. Since $d=O(\log n)$, we conclude that the running time performance of all the data structures in Section 5 becomes $\tilde{O}(n^{O(\epsilon)})$. Then, by scaling $\epsilon$, we assume that the running time of these data structures is $\tilde{O}(n^{\epsilon})$. For simplicity, we consider a parameter $p_{\epsilon}$ which is a polynomial of $(1/\epsilon)$ and is an upper bound on the approximation guarantees of all data structures in Section 5. So, we consider the following throughout the entire Section 6.
@@
[…]
@@
Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $\Delta]^{d}$. We call this sequence $t$-robust w.r.t. the current data set $X$, if there exists a sequence $(B_{0},B_{1},\cdots,B_{t})$ of subsets of $X$ such that the following hold,
@@
[…]
@@ 6.2 Implementation
1: for $i=0$ to $\log_{2}k$ do
2:   Using restricted $k$-means (Lemma 5.8), compute a subset $\widehat{S}_{i}\subseteq S_{\textnormal{{init}}}$ of $(k-s_{i})$ centers that is a $p_{\epsilon}$-approximation to $\mathrm{OPT}_{k-s_{i}}^{S_{\textnormal{{init}}}}\left(X^{(0)}\right)$.
3:   if $\textsc{cost}(X^{(0)},\widehat{S}_{i})>p_{\epsilon}^{96}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})$ then
4:    return $\widehat{\ell}:=s_{i-1}$.
After finding $\widehat{\ell}$, we set the length of the epoch to be $\ell+1$ where $\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot p_{\epsilon}^{102}}\right\rfloor$.
@@ 6.3 Implementing Robustify
[…]
@@ 6.4 Searching for Non-Robust Centers
The proof of this claim follows from [BCF24], but for the sake of completeness, we provide the proof in Part III (see Section 21.10).
@@ 6.4.1 Implementing Part 1
This claim shows that every update in the data set can contaminate at most one center in each scale $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$. The proof of this claim follows from [BCF24], but for the sake of completeness, we provide the proof in Part III (see Section 21.11).
@@ 6.4.2 Implementing Part 2
This lemma follows from combining the ANN and [BhattacharyaGJQ24, Theorem 3.1]. We have already implemented ANN in Section 5, and in [BhattacharyaGJQ24], it is explained how we can implement the above data structure by using an ANN.
@@ 6.4.3 Termination of Robustify
After each change in the center set $S$, the distances of centers $u\in S$ to other centers might change. This means that new centers might be inserted into $\mathcal{Y}$ and they might need a subsequent call to MakeRobust. Similar to [BCF24], it is possible to show that if a call MakeRobust is made on a center in Robustify, it will continue to satisfy Conditions (14) and (10) until the end of the call to the current Robustify, and our procedure will not make another call on that center.
@@
[…]
The proof of this lemma follows from [BCF24] (although we have new definitions), and we provide the proof in Part III for the sake of completeness (see Section 21.9).
@@
1: $W_{1}\leftarrow W^{\prime}\setminus S_{\textnormal{{init}}}$. $\triangleright$ All new centers
2: for each $x\in X^{(0)}\oplus X^{(\ell+1)}$ do
3:   for each $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$ do
4:    Let $u\leftarrow\widehat{\mathrm{dist}}(x,S[i]-x)$.
5:    if $\mathrm{dist}(u,x)\leq p_{\epsilon}^{6i+3}$ then
6:      $W_{2}\leftarrow W_{2}+u$. $\triangleright$ Mark $u$ as a contaminated center
7: $S\leftarrow S_{\textnormal{{init}}}-(S_{\textnormal{{init}}}\setminus W^{\prime})+(W^{\prime}\setminus S_{\textnormal{{init}}})$ $\triangleright$ This line updates data structures
8: for each $u\in W_{1}\cup W_{2}$ do
9:   $\texttt{MakeRobust}(u)$.
10: while $\mathcal{Y}\neq\emptyset$ do $\triangleright$ Before this line, many centers might have been inserted to $\mathcal{Y}$
11:   $u\leftarrowpop$$(\mathcal{Y})$
12:   $t\leftarrow$ Smallest integer satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{20}$
13:   if $t[u]<t$ then
14:    $\texttt{MakeRobust}(u)$.
15: return $S_{\textnormal{{final}}}\leftarrow S$ .
@@ 6.5 Make-Robust
In this section, we introduce the subroutine MakeRobust that given $u\in S$, swaps this center with a close center $v$ which is robust (i.e. satisfies Condition Equation 14 as well as Equation 10). We start by computing the smallest integer $t$ satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{14}$. The goal is to follow the Definition 6.1. We define $x_{t}:=u$ and by iterating over $j=t$ down to 0, we find $x_{j-1}$ and $B_{j}$ satisfying the properties in Definition 6.1. We use Lemma 5.5, and make a query on this data structure for $x=x_{j}$ and $\rho=p_{\epsilon}^{6j}$. It returns a center $c^{\star}\in[\Delta]^{d}$, together with numbers $\widehat{\textsc{cost}}_{c^{\star}}$, $\widehat{\textsc{cost}}_{x_{j}}$, and $b_{x_{j}}$.
@@
1: $t\leftarrow$ Smallest integer satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{14}$,   $x_{t}\leftarrow u$
2: for $j=t$ down to $1$ do
3:   Invoke Lemma 5.5 for $x=x_{j}$ and $\rho=p_{\epsilon}^{6j}$ to get $c^{\star}\in[\Delta]^{d}$, $\widehat{\textsc{cost}}_{x_{j}}$, $\widehat{\textsc{cost}}_{c^{\star}}$ and $b_{x_{j}}$
4:   if $\left(\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\geq p_{\epsilon}^{6j-3}\right)$ or $\left(\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\leq\widehat{\textsc{cost}}_{c^{\star}}\right)$ then
5:    $x_{j-1}\leftarrow x_{j}$
6:   else
7:    $x_{j-1}\leftarrow c^{\star}$
8: $v\leftarrow x_{0}$
9: $S\leftarrow S-u+v$
10: Save $t[v]:=t$ together with $v$
@@
The analysis of the approximation ratio of the algorithm follows from the prior work [BCF24] with deliberate changes in the parameters inside the proofs. There are only two important considerations as follows. First, our new definition of robust centers still satisfy the main properties of robust centers as in [BCF24] (see Section 16.1.1). Second, the analysis of [BCF24] depends on the fact that the integrality gap of the standard LP-relaxation for $k$-median is constant. Since the same fact holds for $k$-means (see Theorem 1.1 in [integrality-gap-k-means]), we can extend the analysis for our purpose. For the sake of completeness, we provide a thorough analysis in Part III.
@@
The recourse analysis of the algorithm also follows from the previous work [BCF24] with deliberate changes. There are only some considerations as follows. First, in augmented $k$-means (Lemma 5.11), instead of finding a set $A$ of size exactly $a$, we find a set of size $\tilde{O}(a)$. This increases the recourse of the algorithm by some $\log$ factors, but the same analysis works as in [BCF24]. Second, we changed the Condition (10) to a more restricted Condition (14), and we might call MakeRobust more often. But, the same analysis as in [BCF24] can be applied here to show that the total amortized recourse of Robustify is still $\tilde{O}(1)$. For the sake of completeness, we provide a thorough analysis in Part III.
@@ 7.1.3 Running Time of Robustify
For each of these centers $u\in\mathcal{Y}$, during 11 to 13, the algorithm decides in $\tilde{O}(1)$ time whether or not a call to MakeRobust will be made on uu. [^7] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 5.6. As a result, the amortized time spend on part 2 is $\tilde{O}(n^{\epsilon})$.
@@ 7.1.4 Running Time of MakeRobust
The main for loop in the implementation consists of at most $O(\log_{p_{\epsilon}^{2}}(\sqrt{d}\Delta))$ many iterations. Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 5.5) used in 3 of Algorithm 3. It is obvious that all other parts of the implementation take at most $\tilde{O}(1)$ time. Hence, the total time spent to perform one call to MakeRobust is at most $\tilde{O}(n^{\epsilon})$.
@@ 7.2.2 Correctness of MakeRobust
In this part of the paper, we provide several dynamic data structures that are useful in subsequent parts. We begin by summarizing the formal statement of these data structures in Section 8. As a crucial foundation of our data structures, we design a time-efficient consistent hashing scheme in Section 9. Then, we elaborate on the construction and analysis of each data structure in Sections 10, 11, 12 and 13.
@@
[…]
@@ 8 Main Data Structures
The following Lemmas 8.1, 8.2, 8.3 and 8.4 are established in Section 10, where we design a unified range query structure for handling a class of “mergeable” query problems. Specifically, Lemmas 8.1, 8.2 and 8.3 address a set of approximate nearest neighbor (ANN) problems, while Lemma 8.4 concerns the $1$-means problem over approximate balls. Here, an “approximate ball” refers to a subset that is sandwiched between two balls with the same center and (up to a constant) roughly the same radii.
@@
The following Lemmas 8.5 and 8.6 address two crucial subproblems – restricted $k$-means and augmented $k$-means – respectively, which play a central role in our dynamic algorithm described in Part III. We present the data structures of Lemmas 8.5 and 8.6 in Sections 12 and 13, respectively, both of which are built on a common data structure that (implicitly) maintains an approximate assignment from the point set $X$ to the center set $S$, as described in Section 11.
@@
$\textsc{cost}(X,S-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
@@
$\textsc{cost}(X,S+A)\penalty 10000\ \leq\penalty 10000\ O(1)\cdot\min_{A^{\prime}\subseteq[\Delta]^{d}:|A^{\prime}|=a}\textsc{cost}(X,S+A^{\prime}).$
@@ 9 Efficient Consistent Hashing
    (efficiency) For every point $x\in[\Delta]^{d}$, both the hash value $\varphi(x)$ and a size-$(\leq\Lambda)$ subset $\Phi_{x}\subseteq\Phi$ satisfying $\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho))$ can be computed in time $\tilde{O}(\Lambda)$.
@@
For any $\Gamma\geq 2\sqrt{2\pi}$, $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, and $\rho>0$, there exists a random hash $\varphi$ on $\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
@@
Our construction of an efficient consistent hash builds on a previous construction from [random-shift], which provides a weaker guarantee of consistency that holds in expectation. In addition, we leverage the special structure of their hash function – based on randomly shifted grids – and show that the set of hash values over any given ball can be computed efficiently (see Item 4 below). We summarize these results in the following lemma.
Lemma 9.3 (Hashing with Weaker Consistency Guarantees [random-shift]).
@@
Items 1, 2 and 3 are explicitly stated in [random-shift]. We now prove Item 4.
We review the construction from [random-shift]: Let $v\sim\operatorname{Unif}([0,r/\sqrt{d}]^{d})$ be a uniformly random real vector. The hash function is defined as $\tilde{\varphi}:x\mapsto\lfloor\frac{x+v}{r/\sqrt{d}}\rfloor$ for $x\in[\Delta]^{d}$. Herein, the integer vector $\lfloor z\rfloor:=(\lfloor z_{1}\rfloor,\dots,\lfloor z_{d}\rfloor)\in\mathbb{Z}^{d}$ coordinatewise rounds down a real vector $z=(z_{1},\dots,z_{d})\in\mathbb{R}^{d}$.
@@
$\widetilde{\varphi}^{-1}(z)$ $\textstyle\penalty 10000\ =\penalty 10000\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\penalty 10000\ =\penalty 10000\ \prod_{j\in[d]}\big[z_{j}\cdot r/\sqrt{d}-v_{j},(z_{j}+1)\cdot r/\sqrt{d}-v_{j}\big),$
$\widetilde{\varphi}^{-1}(z^{\prime})$ $\textstyle\penalty 10000\ =\penalty 10000\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\penalty 10000\ =\penalty 10000\ \prod_{j\in[d]}\big[z^{\prime}_{j}\cdot r/\sqrt{d}-v_{j},(z^{\prime}_{j}+1)\cdot r/\sqrt{d}-v_{j}\big).$
@@
  • $\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\leq\rho$, for every hash value $z\in\mathbb{Z}^{d}$.
    $\triangleright$ Item 1 of Lemma 9.3
  • $\mathbb{E}\big[|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\big]\leq\frac{\Lambda}{2C}$, for every point $x\in[\Delta]^{d}$.
    $\triangleright$ Item 2 of Lemma 9.3
@@
$\varphi(x)\penalty 10000\ :=\penalty 10000\ (c_{x},\widetilde{\varphi}_{c_{x}}(x)),$ $\forall x\in[\Delta]^{d}.$ (17)
@@
First, the diameter bound and the consistency properties (Items 1 and 2 of Definition 9.1) follow from Claims 9.5 and 9.6, respectively.
@@
By construction, our consistent hash function $\varphi(x)=(c_{x},\widetilde{\varphi}_{c_{x}}(x))$ induces a natural partition of the input space $\Delta]^{d}=\cup_{c\in[C]}G_{c}$, where $G_{c}:=\{x\in[\Delta]^{d}\mid c_{x}=c\}$, $\forall c\in[C]$. In this manner, we have $|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\sum_{c\in[C]}|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c})|$, so it suffices to show that
@@
Second, the efficiency guarantee (Item 3 of Definition 9.1) directly follows from a combination of Claims 9.7 and 9.8.
@@
Following Lemma 9.3 and the stop condition $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, the depth-first search for every single $c=1,2,\dots,C$ stops in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$.
@@
Following Lemma 9.3 and the stop condition $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, every iteration $c=1,2,\dots,C$ will stop in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$ and augment the subset $\Phi_{x}$ by at most $\frac{\Lambda}{C}$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$ and the total size $|\Phi_{x}|\leq\Lambda$.
@@
(ii) $z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$, given our construction of the hash function $\varphi$ (Equations 16 and 17).
@@
$\mathrm{dist}(x,y)\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,\widetilde{\varphi}_{c}^{-1}(z))+\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\penalty 10000\ \leq\penalty 10000\ \tfrac{\rho}{\Gamma}+\rho\penalty 10000\ \leq\penalty 10000\ 2\rho.$